# Analysis of a mathematical model for malaria using data-driven approach

### v1

@@ 2.3 Temperature and altitude dependence of the transmission rate
Thus we observe that the transmission rate is the highest when the height is $150$ and in this case, it is the highest value among the chosen values thus from this we can conclude that $h=150$ is the threshold value for the altitude. The function $f(h)=e^{-\frac{h^{2}}{\xi^{2}}}(1-e^{-\frac{h^{2}}{\xi^{2}}})$ achieves its highest value when $h=\xi{\sqrt{\ln(2)}}$ and thus the transmission rate increases as the value of height approaches $h={\xi\sqrt{\ln(2)}}$.
By substituting both the values of $\xi$ and taking the average we get approximately $142.1$ thus the optimal value for altitude can be concluded to be in the range $140-150$ and since it can be observed that the trajectory for all of these values is nearly the same.
3 Numerical Simulation and Deep Learning
Whenever there is a disease transmission in a particular country or state, there is a significant number of deaths and sufferings that the people of the place have to undergo. One of the main reasons why this happens is when the spread of the disease goes uncontrollably. In this work, the following problem statements are addressed to understand the transmission of disease.
3.1 Predicting the trajectory of the compartments
Whenever there is a disease transmission in a country/state the medical professionals and the government professionals should be prepared to handle the situation and to provide different levels of support such as vaccination, hospital facilities, implementing curfews, etc. Knowing the trajectory will enable the government to make more precise decisions. The overall methodology for this problem statement can be found in Figure 27.
Here the neural network is used to make a recursive relation between the trajectory at $t=n$ and trajectory $t=n+1$. The main reason why neural network is used is because it has a unique ability to identify nonlinear relations in the data.
Following are the parameters that are used for predicting the trajectory:
$\beta_{h}$ $\beta_{m}$ $\gamma$ $\mu_{h}$ $\Gamma_{h}$ $\mu_{m}$ $\Gamma_{m}$
0.9 0.8 0.05 0.02 0.02 0.2 0.08
Figure 27: Neural network architecture and methodology
Here the neural network is used to make a recursive relation between the trajectory at $t=n$ and trajectory $t=n+1$. Neural network is mainly used because of its unique ability to identify nonlinear relations in the data. The predicted and the actual trajectories are shown in Figure 28 - Figure 29.
(a) Susceptible human population
(b) Infected human population
Figure 28: Comparision of predicted and actual trajectories for susceptible and infected human population.
(a) Susceptible mosquito population
(b) Infected mosquito populatio
Figure 29: Comparision of predicted and actual trajectories for susceptible and infected mosquito population.
Now every compartment has two factors one is called the incoming factor which means the factors that lead to an increase in the compartment’s population. The other one is the outgoing factor which leads to a decrease in the compartment’s population.
In the susceptible compartment, there is only one incoming factor which is the birth rate, and where there are two outgoing factors namely the death rate and the interaction with the infected population this can be easily observed from . Since the number of outgoing factors is more than the number of incoming factors the susceptible population decreases with time.
Similar to the susceptible compartment the infected compartment also has only one incoming factor which is the interaction between the susceptible population and two outgoing factors which are the death rate and the recovery rate. Thus when an infection spreads in a place the number of infected people increases but as time passes the people develop immunity to the disease and thus the infected population decreases after reaching a threshold value.
From the flow chart of the compartment model, we can observe that the transfer of people is from susceptible to infected. Thus we can observe a sudden drop in the susceptible people and a huge rise in the infected compartment at the same time.
3.2 PREDICTING THE TRANSMISSION RATE OF THE DISEASE
The main factors that control the dynamics of a disease system are the parameters associated with it. Thus knowing the parameters of disease will enable the medical professionals to know the movement of people from one compartment to the other and it will provide them with a good clarity of what measures must be taken. To address this problem statement three different architectures of neural networks are used in this work they are ANN, CNN, and RNN. The complete overview of the problem can be seen in Figure 30
Figure 30: Details of neural network architecture and methodology
The number of layers for each model is kept the same for a novel comparison. Here the main reason why one trajectory is used for testing is because every disease transmission is always associated with one trajectory and one set of parameters. The results from each of the neural networks can be found in  34,  5 .
Parameters Predicted value Actual value Error
Birth rate of humans 0.056 0.044 28.6 %
Transmission rate of humans 0.227 0.264 14.07 %
Death rate of human 0.442 0.451 1.86 %
Recovery rate of humans 0.378 0.379 0.32 %
Birth rate of mosquitoes 0.450 0.444 1.37 %
Transmission rate in mosquitoes 0.242 0.241 0.40 %
Death rate of mosquitoes 0.3806 0.3807 0.04 %
Table 3: Parameter prediction using ANN
Parameters Predicted value Actual value Error
Birth rate of humans 0.1904 0.1784 6.69 %
Transmission rate of humans 0.264 0.297 $1$
Death rate of human 0.428 0.433 1.02 %
Recovery rate of humans 0.1605 0.1602 0.17 %
Birth rate of mosquitoes 0.223 0.217 2.96 %
Transmission rate of mosquitoes 0.284 0.274 3.54 %
Death rate of mosquitoes 0.479 0.481 0.24 %
Table 4: Parameter prediction using CNN
Parameters Predicted value Actual value Error
Birth rate of humans 0.172 0.169 1.44 %
Transmission rate of humans 0.111 0.103 8.21 %
Death rate of human 0.0854 0.0841 1.58 %
Recovery rate of humans 0.492 0.495 0.17 %
Birth rate of mosquitoes 0.445 0.457 2.47 %
Transmission rate of mosquitoes 0.334 0.338 1.15 %
Death rate of mosquitoes 0.464 0.482 3.38 %
Table 5: Parameter prediction using RNN
For each model, the epoch was set to 10000 to generate the corresponding result and from the comparison, we can conclude that RNN’s performance was the best of all the three. The main reason RNN’s performance is the best is because RNN can store large amounts of data and make accurate predictions from it compared to the other two neural networks.
@@ 3.3 FINDING THE RISK OF A DISEASE
The most important aspect of a disease is the determination of risk. Whenever there is a disease outbreak in a country there are some regions where there is more risk compared to the other regions thus it is essential to calculate this risk of every region. This problem statement is addressed using the method of DMD ( dynamic mode decomposition ) and the main reason for using this method is because DMD can make exact predictions from raw data, unlike other deep learning methods. The complete methodology of the problem statement can be seen in Figure  31.
In this work, DMD is used to calculate the disease risk in a particular region.DMD gives the oscillations of the dynamics and thus using the peak values of DMD will give us the overall oscillations of the dynamics which is nothing but the measure of risk. The DMD plot and the eigenvalue spectrum can be found in Figure  32.
From the eigenvalue spectrum, we can observe that all the points are either on or within the unit circle. This shows that the transmission of malaria in Africa is stable. African map with the corresponding color coding based on the risk can be found in Figure  33.
@@ 4 Concluding Remarks
In this work, an attempt is made to understand the dynamics of malaria transmission using various mathematical and machine-learning techniques. Using neural networks, the trajectories of all five compartments are predicted, followed by predicting the parameters of disease dynamics using three different architectures of neural networks and finally, this work was concluded by finding the measure of risk using DMD. One thing that could be understood is that disease transmissions are not random processes, but happen under the influence of underlying physical laws.
This work can be deployed by the government of any country to analyze the trajectory of infected people whenever there is a disease outbreak and then predict the future state of the country which will give medical professionals an idea of what kind of measures must be taken to reduce the disease spread
In the future, our idea is to understand the dynamics of malaria using techniques of physics-informed machine learning such as Physics neural networks (PINNs) and Sparse identification of non-linear dynamics (SINDy). We also want to study the stochastic version of this model mainly to study the random fluctuations in the birth and mortality process.

### v2

@@ 2.4 Temperature and altitude dependence of the transmission rate
Thus we observe that the transmission rate is the highest when the height is $150$ and in this case, it is the highest value among the chosen values thus from this we can conclude that $h=150$ is the threshold value for the altitude. The function $f(h)=e^{-\frac{h^{2}}{\xi^{2}}}(1-e^{-\frac{h^{2}}{\xi^{2}}})$ achieves its highest value when $h=\xi{\sqrt{\ln(2)}}$ and thus the transmission rate increases as the value of height approaches $h={\xi\sqrt{\ln(2)}}$. By substituting both the values of $\xi$ and taking the average we get approximately $142.1$ thus the optimal value for altitude can be concluded to be in the range $140-150$ and since it can be observed that the trajectory for all of these values is nearly the same.
3 Parameter estimation
An important factor influencing disease transmission is the parameters associated with the transmission dynamics. Understanding these parameters is crucial for medical professionals, as it enables them to assess the severity of the disease and implement appropriate interventions. This section presents three distinct neural network models: Artificial Neural Networks (ANNs), Recurrent Neural Networks (RNNs), and Physics-Informed Neural Networks (PINNs) for estimating these parameters. Utilizing the predicted parameters, we can forecast the trajectories of the various compartments involved in the disease dynamics.
For this study, the Artificial Neural Networks (ANNs) utilized comprised a total of five layers, including three hidden layers. All layers, except for the output layer, contained 15 dense units and employed the sigmoid function as the activation function. The output layer featured seven dense units with no activation function applied. The Recurrent Neural Networks (RNNs) architecture implemented in this work consisted of three layers: one input layer, one dropout layer, and one output layer. The input layer was composed of 50 LSTM units with the ReLU activation function. The dropout layer had a dropout rate of 20 %, while the output layer contained seven dense units without an activation function. The Physics-Informed Neural Networks (PINNs) architecture mirrored that of the ANN, with the primary distinction being the number of nodes in the input and output layers, which were one and five, respectively, in the PINNs architecture.
3.1 Methodology
Since three different architectures of neural networks are used there is a difference between methodologies used for ANN, RNN, and PINN. For ANN and RNN, both of these models have been trained on a training dataset that has 1000 data points where the input is the first 10 points of the trajectories of all of the compartments and the output is the corresponding parameters. For PINN the methodology is completely different from the earlier models since here there is no involvement of a training dataset but instead an assumed set of parameters is chosen and by minimizing the loss function the assumed set of parameters is made to converge to the actual set of parameters by comparing the obtained and actual trajectories. Here the input is the time whereas the output is the trajectories of all the five compartments.
For a novel comparison between the models, each model is simulated for 20000 epochs. Details of actual and predicted parameters are given in the tables below.
Parameter Predicted Value Actual Value Relative Error (%)
Birth rate of humans 0.38273 0.40000 4.32
Transmission rate of humans 0.28249 0.30000 5.84
Death rate of humans 0.09973 0.10000 0.27
Recovery rate of humans 0.04640 0.01000 364.01
Birth rate of mosquitoes 0.07613 0.05000 52.25
Transmission rate of mosquitoes 0.01045 0.02000 47.73
Death rate of mosquitoes 0.05757 0.04000 43.92
Table 3: Comparison of predicted and actual values with relative errors for ANN
Parameter Predicted Value Actual Value Relative Error (%)
Birth rate of humans 0.19740 0.40000 50.65
Transmission rate of humans 0.18619 0.30000 37.94
Death rate of humans 0.34348 0.10000 243.48
Recovery rate of humans 0.21954 0.01000 2095.38
Birth rate of mosquitoes 0.18898 0.05000 277.96
Transmission rate of mosquitoes 0.24995 0.02000 1149.77
Death rate of mosquitoes 0.38605 0.04000 865.13
Table 4: Comparison of predicted and actual values with relative errors for RNN
Parameter Predicted Value Actual Value Relative Error (%)
Birth rate of humans 0.39200 0.40000 2.00
Transmission rate of humans 0.28186 0.30000 6.05
Death rate of humans 0.09793 0.10000 2.07
Recovery rate of humans 0.00949 0.01000 5.07
Birth rate of mosquitoes 0.04593 0.05000 8.14
Transmission rate of mosquitoes 0.01969 0.02000 1.56
Death rate of mosquitoes 0.03600 0.04000 10.01
Table 5: Comparison of predicted and actual values with relative errors for PINN
3.2 Prediction of trajectories
We observe that out of all the models PINNs can be observed to be superior to the remaining models. One of the main reasons why PINNs proved to be a superior model is because disease transmissions are not random phenomena but are phenomena following a certain law and thus PINNs not only model the data but also model the associated physics law thus enabling the models to make predictions with very high accuracy. The parameters predicted by PINNs and the trajectories of all five compartments are predicted. The evolution of the human population can be observed in Figure 28 and the evolution of the mosquito population can be observed in Figure 29.
Figure 28: Prediction made by PINNs for the human population.
Figure 29: Prediction made by PINNs for the mosquito population.
@@ 3.3 Finding the risk of a disease
The most important aspect of a disease is the determination of risk. Whenever there is a disease outbreak in a country there are some regions where there is more risk compared to the other regions thus it is essential to calculate this risk of every region. This problem statement is addressed using the method of DMD (dynamic mode decomposition) and the main reason for using this method is because DMD can make exact predictions from raw data, unlike other deep learning methods. The complete methodology of the problem statement can be seen in Figure  30.
In this work, DMD is used to calculate the disease risk in a particular region. DMD gives the oscillations of the dynamics and thus using the peak values of DMD will give us the overall oscillations of the dynamics which is nothing but the measure of risk. The DMD plot and the eigenvalue spectrum can be found in Figure  31.
From the eigenvalue spectrum, we can observe that all the points are either on or within the unit circle. This shows that the transmission of malaria in Africa is stable. African map with the corresponding color coding based on the risk can be found in Figure  32.
@@ 4 Concluding remarks
In this work, an attempt is made to understand the dynamics of malaria transmission using various mathematical and machine-learning techniques. The proposed theorems related to steady states were validated using numerical simulations it was followed by Parameter estimation using three different architectures of neural networks namely ANNs, RNNs, and PINNs, and the predicted parameter of the best model was then utilized to predict the trajectories of all the compartments. Finally, this work was concluded by finding the measure of risk using DMD.
This mathematical model developed in this work can be deployed by the government of any country to analyze the trajectory of infected people whenever there is a disease outbreak and then predict the future state of the country which will give medical professionals an idea of what kind of measures must be taken to reduce the disease spread
In the future, our idea is to understand the dynamics of malaria using techniques of physics-informed machine learning such as Sparse identification of non-linear dynamics (SINDy), Universal differential equations(UDE), Neural differential equations (NeuralODE), and other variants of PINN’s like VPINNs and Recurrent PINNs. We intend to predict the trajectories of the compartments using time series models such as Attention models, ResNet architecture, etc using a publicly available dataset for epidemic forecasting. We also want to study the stochastic version of this model mainly to study the random fluctuations in the birth and mortality process.